OpenStudy (cucacula):

find dy/dx of the following : 6xy^3 - x^2 = 8 anyone?

sam (.sam.):

Do you know how to start? this is implicit differentiation, so, for 6xy^3 use chain rule for -x^2 power rule, and differentiation of 8 is 0.

sam (.sam.):

for 6xy^3 use it with product rule

sam (.sam.):

\[6xy^3 - x^2 = 8 \\ \\ 6[x(3)(y^2)(y')+y^3]-2x=0\] \[18xy^2y'=2x \\ \\ y'=\frac{2x}{18xy^2} \\ \\y'=\frac{1}{9y^2}\]

OpenStudy (cucacula):

how you get (y^1) + y^3? @.Sam.

sam (.sam.):

Mistake strikes again :| , should be \[6xy^3 - x^2 = 8 \\ \\ 6[x(3)(y^2)(y')+y^3]-2x=0\] \[18xy^2y'=2x-6y^3 \\ \\ y'=\frac{2x-6y^3}{18xy^2} \\ \\y'=\frac{2(x-3y^3)}{18xy^2} \\ \\y'=\frac{x-3y^3}{9xy^2}\]

sam (.sam.):

that's not y^1, that's y "Primed", means dy/dx

sam (.sam.):

\[\Huge y'=\frac{dy}{dx} \\ \\ \Huge y^1 \neq \frac{dy}{dx}\]

sam (.sam.):

You get (y^1) + y^3 from product rule \[\Large \frac{d}{dx}(uv)=u\frac{dv}{dx} + v\frac{du}{dx} \]

OpenStudy (cucacula):

need to find U and V again right? @.Sam.

sam (.sam.):

Noneed, from 6xy^3, "u" is the x, "v" is the y^3 so using product rule, leave the 6 out, \[6[x(3)y^2(\frac{dy}{dx})+y^3(1)]\]

sam (.sam.):

Everytime you differentiate "y" you will get dy/dx

OpenStudy (cucacula):

I know already but im still confuse with the y^3(1) @.Sam.

sam (.sam.):

That part is \[v\frac{du}{dx}\] From the product rule. We said that "v" is the y^3 , and for du/dx means the differentiation of "u", differentiate "x" is 1, so we get y^3(1) or just y^3

OpenStudy (cucacula):

thank you @.Sam. for helping Im already solved it :)

sam (.sam.):

np :)